### v1

@@ Abstract
In this study, we introduce a methodology for human image animation by leveraging a 3D human parametric model within a latent diffusion framework to enhance shape alignment and motion guidance in curernt human generative techniques. The methodology utilizes the SMPL(Skinned Multi-Person Linear) model as the 3D human parametric model to establish a unified representation of body shape and pose. This facilitates the accurate capture of intricate human geometry and motion characteristics from source videos. Specifically, we incorporate rendered depth images, normal maps, and semantic maps obtained from SMPL sequences, alongside skeleton-based motion guidance, to enrich the conditions to the latent diffusion model with comprehensive 3D shape and detailed pose attributes. A multi-layer motion fusion module, integrating self-attention mechanisms, is employed to fuse the shape and motion latent representations in the spatial domain. By representing the 3D human parametric model as the motion guidance, we can perform parametric shape alignment of the human body between the reference image and the source video motion. Experimental evaluations conducted on benchmark datasets demonstrate the methodology’s superior ability to generate high-quality human animations that accurately capture both pose and shape variations. Furthermore, our approach also exhibits superior generalization capabilities on the proposed wild dataset. Project page: https://fudan-generative-vision.github.io/champ.
@@
footnotetext: $*{}^{*}$ These authors contributed equally to this work.footnotetext: ${}^{\dagger}$ Corresponding Author
@@ 1 Introduction
Recent advancements in generative diffusion models, particularly latent diffusion models, have significantly propelled the field of image animation forward [51, 53, 56]. These advancements have found broad application in virtual reality experiences, interactive storytelling, and digital content creation, resulting in the production of a plethora of sophisticated dynamic visual content. Within the realm of human image animation, techniques typically rely on a reference image and motion guidance specific to humans, such as skeletons [43, 16], semantic maps [36, 47], and dense motion flows [56, 18], to generate controllable human animation videos. In this domain, two predominant approaches prevail: those based on GANs [9, 24] and diffusion models [14, 39].
GAN based methods [35, 44, 40, 45] commonly employ warping functions to spatially transform the reference image according to input motion for generating sequential video frames. By leveraging the inherent generative visual capabilities of GANs, these methods aim to fill in missing regions and improve visually implausible areas within the generated content. Despite yielding promising results in dynamic visual content generation, GAN-based approaches often encounter challenges in effectively transferring motion, particularly in scenarios involving substantial variations in human identity and scene dynamics between the reference image and the source video motion. This limitation manifests as unrealistic visual artifacts and temporal inconsistencies in the synthesized content.
@@
Figure 2: The overview of our proposed approach.
Incorporating the SMPL model as a guiding framework for both shape and pose within a latent diffusion model, our methodology is structured around three fundamental components: 1) The sequences derived from the SMPL model corresponding to the source video are projected onto the image space, resulting in the generation of depth images, normal maps, and semantic maps that encapsulate essential 3D information. The depth image plays a critical role in capturing the 3D structure of the human form, while the normal map enables the representation of orientation and surface details of the human body. The semantic map aids in accurately managing interactions between different components of the human body during the process of animation generation. 2) Our analysis demonstrates that incorporating skeleton-based motion guidance enhances the precision of guidance information, particularly for intricate movements such as facial expressions and finger movements. As a result, the skeleton is maintained as an auxiliary input to complement the aforementioned maps. 3) In the process of integrating depth, normal, semantic, and skeleton maps through feature encoding, the employment of self-attention mechanisms facilitates the feature maps, processed via self-attention, in learning the representative saliency regions within their respective layers. Such multi-layer semantic fusion enhances the model’s capability to comprehend and generate human postures and shapes. Finally, the inclusion of these multi-layer feature embeddings conditioned on a latent video diffusion model leads to precise image animation both in pose and shape.
@@ 2 Related Work
Diffusion Models for Human Image Animation. The task of animating human images, a significant endeavor within the domain of video generation, aims to seamlessly create videos from one or multiple static images [6, 30, 35, 36, 51, 54, 56, 50]. The recent advancements of diffusion models in the text-to-image domain have sparked interest in exploring their utility for animating human images. PIDM [4] introduces a texture diffusion module that is specifically crafted to align the texture patterns of the source and target images closely, thereby enhancing the realism of the resultant animated output. DreamPose [18] capitalizes on the capabilities of the pre-trained Stable Diffusion model by incorporating both CLIP [28] and VAE [19] for image encoding. It integrates these embeddings with an adapter. Similarly, DisCo [43] innovatively segregates the control of pose and background using dual independent ControlNets [53], providing finer control over the animation process. Animate Anyone [16] utilizes a UNet-based ReferenceNet to extract features from reference images. It includes pose information via a lightweight pose guider. Expanding on the principles introduced by AnimateDiff [11], Animate Anyone integrates a temporal layer into the denoising UNet to enhance temporal coherence. MagicAnimate [47] follows a similar approach but employs a ControlNet tailored for DensePose [10] inputs instead of the more commonly used OpenPose [5] keypoints to provide more precise pose guidance. This paper primarily builds upon esteemed diffusion-based methodologies and advances the optimization of appearance alignment and motion guidance mechanisms. This is achieved by introducing a 3D parametric model for geometric reconstruction of the reference image and motion modeling of the source video sequence.
@@ 3.2 Multi-Layer Motion Condition
SMPL to Guidance Conditions. Given a reference human image $I_{\text{ref}}$ and a sequence of driving motion video frames $I^{1:N}$, where $N$ denotes the total number of frames, we obtain the 3D human parametric SMPL model, $H_{\text{ref}}$ and $H_{\text{m}}^{1:N}$, respectively, utilizing an existing framework known as 4D-Humans [8]. In order to extract comprehensive visual information from the pixel space, we render the SMPL mesh to obtain 2D representations. This includes encoding depth maps, which contain distance information from each pixel to the camera, crucial for reconstructing the 3D structure of the scene. Additionally, we encode normal maps, which depict the surface orientation at each point in the image and can capture detailed information about the human body surface. Furthermore, semantic segmentation maps provide class information for each pixel in the image, enabling accurate handling of interactions between different components of the human body.
@@ 3.3 Multi-Layer Motion Guidance
Guidance Self-Attention. ControlNet [53] is frequently used in human animation tasks to control generated actions considering additional guidance. However, introducing multiple guidance condition to ControlNet would result in a computational burden that is unaffordable. In light of this, we are inspired by the advanced work [16] and propose a guidance encoder designed to encode our multilevel guidance. Through this approach, we achieve the simultaneous extraction of information from the guidance while fine-tuning a pre-trained denoising U-Net. The encoder consists of a series of lightweight networks. We assign a guidance network $\mathcal{F}^{i}(\cdot,\theta^{i})$ to each guidance condition to encode its features. For each guidance network, we first extract features of the guidance condition through a set of convolutional layers. Considering the presence of multilevel guidance conditions, which involve different characteristics of the human body, a self-attention module is appended after the convolutional layers. This module facilitates the precise capture of corresponding semantic information for each of the multi-layer guidance condition. In particular, Figure 3 illustrates the self-attention map of depth, normal, semantic, and skeleton feature embeddings post-training. The analysis reveals distinct patterns: the depth condition predominantly focuses on the geometric contours of the human figure; the normal condition emphasizes the orientation of the human body; the semantic condition prioritizes the semantic information of different body parts; and the skeleton attention provides detailed constraints on the face and hands.
Multi-Layer Motion Fusion. In order to preserve the integrity of the pretrained denoising U-Net model, we opt to apply zero convolution to extract the features of each guidance condition. The guidance encoder consolidates the feature embeddings from all the guidance conditions by aggregating them through summation, yielding the ultimate guidance feature denoted as $y$. This operation can be expressed mathematically as:
@@
where $N$ signifies the total count of guidance conditions incorporated. Subsequently, the guidance feature is combined with the noisy latent representation before being fed into the denoising fusion module.
@@ 3.4 Network
Inference. During the inference process, animation is performed on a specific reference image by aligning the motion sequences extracted from wild videos or synthesized ones. Parametric shape alignment is utilized to align the motion sequence with the reconstructed SMPL model derived from the reference image at the pixel level, providing a basis for animation. To accommodate the input of a video clip comprising 24 frames, a temporal aggregation technique [41] is employed to concatenate multiple clips. This aggregation method aims to produce a long-duration video output.
@@ 4.1 Implementations
Dataset. We have curated a dataset comprising approximately 5,000 high-fidelity authentic human videos sourced from reputable online repositories, encompassing a total of 1 million frames. The dataset is segmented as follows: Bilibili (2540 videos), Kuaishou (920 videos), Tiktok & Youtube (1438 videos), and Xiaohongshu (430 videos). These videos feature individuals of varying ages, ethnicities, and genders, depicted in full-body, half-body, and close-up shots, set against diverse indoor and outdoor backdrops. In order to enhance our model’s capacity to analyze a wide range of human movements and attire, we have included footage of dancers showcasing various dance forms in diverse clothing styles. In contrast to existing datasets characterized by pristine backgrounds, our dataset capitalizes on the diversity and complexity of backgrounds to aid our model in effectively distinguishing the foreground human subjects from their surroundings. To maintain fairness and align with established benchmarks in the field of image animation, the identical test set as utilized in MagicAnimate [47] has been employed for TikTok evaluation.
@@
Table 1: Quantitative comparisons on Tiktok dataset. * indicates that the proposed approach is fine-tuned on the Tiktok training data-set.
@@ 4.2 Comparisons
Figure 4: Qualitative comparisons between our and the state-of-the-art approaches on the benchmark data-sets.
@@
Evaluation on TikTok dataset. Table 1 presents a concise quantitative analysis of various methods evaluated on the TikTok dataset, focusing on key metrics such as L1 loss, PSNR, SSIM, LPIPS, FID-VID, and FVD. The proposed method, both in its original and fine-tuned (* indicated) forms, demonstrates superior performance across most metrics, particularly highlighting its effectiveness in achieving lower L1 loss, higher PSNR and SSIM values, and reduced LPIPS, FID-VID, and FVD scores. Notably, the fine-tuned version of our approach shows the best overall results, indicating the benefits of dataset-specific optimization. Figure 4 provides additional qualitative comparison on such benchmark.
Evaluation on Proposed Unseen Dataset. In order to further compare the robustness of various methods, distinct from datasets such as TikTok and UBC fashion that exhibit domain proximity, we have constructed a testing dataset comprising 100 high-fidelity authentic human videos sourced from online repositories. These videos exhibit significant variations in the shape, pose, and appearance of the individuals depicted. Figure 5 provides some qualitative comparison of the unseen data-sets along with the statistical comparison presented in Table 2, collectively illustrate the efficacy of the proposed approach in generalizing to unseen domains.
@@ 4.3 Ablation Studies
Different Conditions from SMPL. As shown in Table 3, the statistics demonstrate that the full configuration of the proposed method (“ours”) obviously outperforms other variants in terms of image quality and fidelity, and video consistency and realism. The ablation components, “w/o SMPL”, “w/o geometry”, and “w/o skeleton”, show progressively improved performance as more components are included. Specifically, SMPL brings obviously more gains in PSNR (1.27 v.s. 0.48) and SSIM (0.10 v.s. 0.05) gains than skeleton, which means better preserve the shape alignment and motion guidance. Moreover, it is noteworthy that the incorporation of both the SMPL and skeleton models leads to additional improvements. Specifically, the skeleton model exhibits advantages in providing refined motion guidance for facial and hand regions. Meanwhile, Figure 9 qualitative demonstrates the effectiveness of different conditions from SMPL.
@@ 4.4 Limitations and Future Works
Despite the enhanced shape alignment capabilities and motion guidance offered by the human parametric SMPL model, its modeling capacity for faces and hands is limited. Consequently, the guidance effect for faces and hands does not match the efficacy of feature-based methods. This limitation underpins the incorporation of DWpose as an additional constraint for facial and hand modeling. As illustrated in Figure 3, the mechanism of self-attention further amplifies the saliency of faces and hands within the skeleton map. However, it is important to note that since SMPL and DWpose are solved independently, a potential discrepancy in consistency between them exists. Although this discrepancy did not manifest significantly in our experiments, it nonetheless introduces a potential source of error.

### v2

@@ Abstract
In this study, we introduce a methodology for human image animation by leveraging a 3D human parametric model within a latent diffusion framework to enhance shape alignment and motion guidance in curernt human generative techniques. The methodology utilizes the SMPL(Skinned Multi-Person Linear) model as the 3D human parametric model to establish a unified representation of body shape and pose. This facilitates the accurate capture of intricate human geometry and motion characteristics from source videos. Specifically, we incorporate rendered depth images, normal maps, and semantic maps obtained from SMPL sequences, alongside skeleton-based motion guidance, to enrich the conditions to the latent diffusion model with comprehensive 3D shape and detailed pose attributes. A multi-layer motion fusion module, integrating self-attention mechanisms, is employed to fuse the shape and motion latent representations in the spatial domain. By representing the 3D human parametric model as the motion guidance, we can perform parametric shape alignment of the human body between the reference image and the source video motion. Experimental evaluations conducted on benchmark datasets demonstrate the methodology’s superior ability to generate high-quality human animations that accurately capture both pose and shape variations. Furthermore, our approach also exhibits superior generalization capabilities on the proposed in-the-wild dataset. Project page: https://fudan-generative-vision.github.io/champ.
@@
footnotetext: These authors contributed equally to this work.footnotetext: Corresponding Author
@@ 1 Introduction
Recent advancements in generative diffusion models, particularly latent diffusion models, have significantly propelled the field of image animation forward [61, 63, 66, 25]. These advancements have found broad application in virtual reality experiences, interactive storytelling, and digital content creation, resulting in the production of a plethora of sophisticated dynamic visual content. Within the realm of human image animation, techniques typically rely on a reference image and motion guidance specific to humans, such as skeletons [52, 19], semantic maps [45, 56], and dense motion flows [66, 23], to generate controllable human animation videos. In this domain, two predominant approaches prevail: those based on GANs [12, 30] and diffusion models [17, 48].
GAN based methods [44, 53, 49, 54, 59, 42] commonly employ warping functions to spatially transform the reference image according to input motion for generating sequential video frames. By leveraging the inherent generative visual capabilities of GANs, these methods aim to fill in missing regions and improve visually implausible areas within the generated content. Despite yielding promising results in dynamic visual content generation, GAN-based approaches often encounter challenges in effectively transferring motion, particularly in scenarios involving substantial variations in human identity and scene dynamics between the reference image and the source video motion. This limitation manifests as unrealistic visual artifacts and temporal inconsistencies in the synthesized content.
@@
Incorporating the SMPL model as a guiding framework for both shape and pose within a latent diffusion model, our methodology is structured around three fundamental components: 1) The sequences derived from the SMPL model corresponding to the source video are projected onto the image space, resulting in the generation of depth images, normal maps, and semantic maps that encapsulate essential 3D information. The depth image plays a critical role in capturing the 3D structure of the human form, while the normal map enables the representation of orientation of the human body. The semantic map aids in accurately managing interactions between different components of the human body during the process of animation generation. 2) Our analysis demonstrates that incorporating skeleton-based motion guidance enhances the precision of guidance information, particularly for intricate movements such as facial expressions and finger movements. As a result, the skeleton is maintained as an auxiliary input to complement the aforementioned maps. 3) In the process of integrating depth, normal, semantic, and skeleton maps through feature encoding, the employment of self-attention mechanisms facilitates the feature maps, processed via self-attention, in learning the representative saliency regions within their respective layers. Such multi-layer semantic fusion enhances the model’s capability to comprehend and generate human postures and shapes. Finally, the inclusion of these multi-layer feature embeddings conditioned on a latent video diffusion model leads to precise image animation both in pose and shape.
@@ 2 Related Work
Diffusion Models for Human Image Animation. The task of animating human images, a significant endeavor within the domain of video generation, aims to seamlessly create videos from one or multiple static images [8, 37, 44, 45, 61, 64, 66, 60, 43, 20, 1, 6, 34, 10, 22]. The recent advancements of diffusion models in the text-to-image domain have sparked interest in exploring their utility for animating human images. PIDM [5] introduces a texture diffusion module that is specifically crafted to align the texture patterns of the source and target images closely, thereby enhancing the realism of the resultant animated output. DreamPose [23] capitalizes on the capabilities of the pre-trained Stable Diffusion model by incorporating both CLIP [35] and VAE [24] for image encoding. It integrates these embeddings with an adapter. Similarly, DisCo [52] innovatively segregates the control of pose and background using dual independent ControlNets [63], providing finer control over the animation process. Animate Anyone [19] utilizes a UNet-based ReferenceNet to extract features from reference images. It includes pose information via a lightweight pose guider. Expanding on the principles introduced by AnimateDiff [14], Animate Anyone integrates a temporal layer into the denoising UNet to enhance temporal coherence. MagicAnimate [56] follows a similar approach but employs a ControlNet tailored for DensePose [13] inputs instead of the more commonly used OpenPose [7] keypoints to provide more precise pose guidance. This paper primarily builds upon esteemed diffusion-based methodologies and advances the optimization of appearance alignment and motion guidance mechanisms. This is achieved by introducing a 3D parametric model for geometric reconstruction of the reference image and motion modeling of the source video sequence.
@@
Figure 2: The overview of our proposed approach. Given an input human image and a reference video depicting a motion sequence. We obtain the pose sequence corresponding to the reference image through Parametric Shape Alignment as 3D motion guidance. MLMF is employed to encode multi-layer 3D-related motion information. Referencenet and Temporal-attention ensure identity consistency and temporal coherence, respectively.
@@ 3.2 Multi-Layer Motion Condition
SMPL to Guidance Conditions. Given a reference human image $I_{\text{ref}}$ and a sequence of driving motion video frames $I^{1:N}$, where $N$ denotes the total number of frames, we obtain the 3D human parametric SMPL model, $H_{\text{ref}}$ and $H_{\text{m}}^{1:N}$, respectively, utilizing an existing framework known as 4D-Humans [11]. In order to extract comprehensive visual information from the pixel space, we render the SMPL mesh to obtain 2D representations. This includes encoding depth maps, which contain distance information from each pixel to the camera, crucial for reconstructing the 3D structure of the scene. Additionally, we encode normal maps, which depict the surface orientation at each point in the image and can capture geometric information related to the orientation of the human body surface. Furthermore, semantic segmentation maps provide class information for each pixel in the image, enabling accurate handling of interactions between different components of the human body.
@@ 3.3 Multi-Layer Motion Guidance
Guidance Self-Attention. ControlNet [63] is frequently used in human animation tasks to control generated actions considering additional guidance. However, introducing multiple guidance condition to ControlNet would result in a computational burden that is unaffordable. In light of this, we are inspired by the advanced work [19] and propose a guidance encoder designed to encode our multilevel guidance. Through this approach, we achieve the simultaneous extraction of information from the guidance while fine-tuning a pre-trained denoising U-Net. The encoder consists of a series of lightweight networks. We assign a guidance network to each guidance condition to encode its features. For each guidance network, we first extract features of the guidance condition through a set of convolutional layers. Considering the presence of multilevel guidance conditions, which involve different characteristics of the human body, a self-attention module is appended after the convolutional layers. This module facilitates the precise capture of corresponding semantic information for each of the multi-layer guidance condition. In particular, Figure 3 illustrates the self-attention map of depth, normal, semantic, and skeleton feature embeddings post-training. The analysis reveals distinct patterns: the depth condition predominantly focuses on the geometric contours of the human figure; the normal condition emphasizes the orientation of the human body; the semantic condition prioritizes the semantic information of different body parts; and the skeleton attention provides detailed constraints on the face and hands.
Multi-Layer Motion Fusion. In order to preserve the integrity of the pretrained denoising U-Net model, we opt to use a convolutional layer with zero initialization as the output layer to extract the features of each guidance condition. The guidance encoder consolidates the feature embeddings from all the guidance conditions by aggregating them through summation, yielding the ultimate guidance feature denoted as $y$. This operation can be expressed mathematically as:
@@
where $N$ signifies the total count of guidance conditions incorporated, $i$ is the index of the pose guidance, and $\theta$ is the input pose image. Subsequently, the guidance feature is combined with the noisy latent representation before being fed into the denoising fusion module.
@@ 3.4 Network
Inference. During the inference process, animation is performed on a specific reference image by aligning the motion sequences extracted from in-the-wild videos or synthesized ones. Parametric shape alignment is utilized to align the motion sequence with the reconstructed SMPL model derived from the reference image at the pixel level, providing a basis for animation. To accommodate the input of a video clip comprising 24 frames, a temporal aggregation technique [50] is employed to concatenate multiple clips. This aggregation method aims to produce a long-duration video output.
@@ 4.1 Implementations
Dataset. We have curated an in-the-wild dataset comprising approximately 5,000 high-fidelity authentic human videos sourced from reputable online repositories, encompassing a total of 1 million frames. The dataset is segmented as follows: Bilibili (2540 videos), Kuaishou (920 videos), Tiktok & Youtube (1438 videos), and Xiaohongshu (430 videos). These videos feature individuals of varying ages, ethnicities, and genders, depicted in full-body, half-body, and close-up shots, set against diverse indoor and outdoor backdrops. In order to enhance our model’s capacity to analyze a wide range of human movements and attire, we have included footage of dancers showcasing various dance forms in diverse clothing styles. In contrast to existing datasets characterized by pristine backgrounds, our dataset capitalizes on the diversity and complexity of backgrounds to aid our model in effectively distinguishing the foreground human subjects from their surroundings. To maintain fairness and align with established benchmarks in the field of image animation, the identical test set as utilized in MagicAnimate [56] has been employed for TikTok evaluation.
@@
Table 1: Quantitative comparisons on Tiktok dataset. * indicates that the proposed approach is fine-tuned on the Tiktok training dataset.
Figure 4: Qualitative comparisons between our and the state-of-the-art approaches on TikTok and proposed unseen dataset.
Figure 5: Qualitative comparisons between our and the state-of-the-art approaches on UBC fashion video datasets.
Figure 6: More qualitative results of our approach on the proposed unseen dataset.
@@ 4.2 Comparisons
Evaluation on benchmark dataset. Table 1 presents a concise quantitative analysis of various methods evaluated on the TikTok dataset, focusing on key metrics such as L1 loss, PSNR, SSIM, LPIPS, FID-VID, and FVD. The proposed method, both in its original and fine-tuned (* indicated) forms, demonstrates superior performance across most metrics, particularly highlighting its effectiveness in achieving lower L1 loss, higher PSNR and SSIM values, and reduced LPIPS, FID-VID, and FVD scores. Notably, the fine-tuned version of our approach shows the best overall results, indicating the benefits of dataset-specific optimization. Figure 4 and Figure 5 provides additional qualitative comparison on such benchmark.
Evaluation on Proposed Unseen Dataset. In order to further compare the robustness of various methods, distinct from datasets such as TikTok and UBC fashion that exhibit domain proximity, we have constructed a testing dataset comprising 100 high-fidelity authentic human videos sourced from online repositories. These videos exhibit significant variations in the shape, pose, and appearance of the individuals depicted. Figure 6 and Figure 7 provides some qualitative comparison of the unseen dataset along with the statistical comparison presented in Table 2, collectively illustrate the efficacy of the proposed approach in generalizing to unseen domains.
@@
Figure 9: Comparision between SHERF (left) and ours (right).
@@
Multi-view Animation. Although our method may not match the direct rendering of 3D human representations for consistent novel views, we employ a sequence of SMPLs for multi-view consistent motion guidance and utilize generative models to achieve satisfactory multi-view results. As shown in Figure 9, we compare the results of our multi-view animation with a representative single-image to 3D human reconstruction method, SHERF [20].
@@ 4.3 Ablation Studies
Different Conditions from SMPL. As shown in Table 3, the statistics demonstrate that the full configuration of the proposed method (“ours”) obviously outperforms other variants in terms of image quality and fidelity, and video consistency and realism. The ablation components, “w/o SMPL”, “w/o geometry”, and “w/o skeleton”, show progressively improved performance as more components are included. Specifically, SMPL obviously brings more gains in PSNR (1.27 v.s. 0.48) and SSIM (0.10 v.s. 0.05) gains than a skeleton, which means better preserving the shape alignment and motion guidance. Moreover, it is noteworthy that the incorporation of both the SMPL and skeleton models leads to additional improvements. Specifically, the skeleton model exhibits advantages in providing refined motion guidance for facial and hand regions. Meanwhile, Figure 12 qualitative demonstrates the effectiveness of different conditions from SMPL.
@@ 4.4 Limitations and Future Works
Despite the enhanced shape alignment capabilities and motion guidance offered by the human parametric SMPL model, its modeling capacity for faces and hands is limited. Consequently, the guidance effect for faces and hands does not match the efficacy of feature-based methods. This limitation underpins the incorporation of DWpose as an additional constraint for facial and hand modeling. As illustrated in Figure 3, the self-attention mechanism further amplifies the saliency of faces and hands within the skeleton map. However, it is important to note that since SMPL and DWpose are solved independently, a potential discrepancy in consistency between them exists. Although this discrepancy did not manifest significantly in our experiments, it nonetheless introduces a potential source of error.